# Compressing multivariate functions with tree tensor networks

### v1

@@ V Application: Solving Non-linear Fredholm Equations with Tree Tensor Networks
[…]
@@
     Remap Variables: $f_{i}(\mathbf{t})\leftarrow f_{i}(\mathbf{x})$triangleright The mapping is achieved via the index relabelling $x_{i,j}\rightarrow t_{i,j}$.
@@
     Integrate (Sec. V): $f_{i}(\mathbf{x})\leftarrow f_{i}(\mathbf{x},\mathbf{t})\prod_{i=1}^{n}\prod_{j=1}^{L}T_{i,j}(t_{i,j})$triangleright $T_{i,j}(t_{i,j})=\frac{1}{2}\ \forall i,j$.
@@
[…]
@@ Acknowledgements.
We would like to thank Matthew Fishman for helpful discussions and advice and input on software design. We also thank Ryan Anselm and Marc Ritter for useful discussions. The authors are grateful for ongoing support through the Flatiron Institute, a division of the Simons Foundation. The tools and methods described in this paper are contained within the ITensorNumericalAnalysis.jl library [61]: an open-source Julia package built on-top of ITensors.jl [62] and ITensorNetworks.jl [63] for constructing and optimizing tree tensor network representations of continuous functions, with the tree structure freely definable by the user.
@@ VII Supplementary Material
=$\qquad$= $\mathcal{T}_{(j,j^{\prime})}(x_{j}+x_{j^{\prime}})_{\alpha_{1},\alpha_{2},\ldots,\alpha_{z_{j}-2},\beta}\qquad z_{j}>2,$
@@ VII.3 Further TCI Data
In Fig. S1 we present further data comparing the effectiveness of the tensor cross interpolation (TCI) algorithm for learning the trivariate probability density function $f(\mathbf{r})\propto\exp(-(\mathbf{r}-\bm{\mu})^{T}M^{-1}(\mathbf{r}-\bm{\mu}))$ where $\mathbf{r}=(x,y,z)$, $\bm{\mu}=(\mu_{x},\mu_{y},\mu_{z})$ is the mean vector, and $M$ is a covariance matrix. Here we sample $M$ from the Lewandowski-Kurowicka-Joe (LKJ) [51] distribution with shape parameter $\eta=1$, which is equivalent to sampling uniformly from the space of all covariance matrices. As in the main text, we find the comb tree systematically outperforms the tensor trains for a given covariance matrix $M$. This is most transparent in Fig. S1c), where the TCI algorithm converges to a value for the infinity norm $\epsilon_{\infty}$ that is orders of magnitude lower for the comb tree than for the tensor trains - despite the train ansatzes utilizing a higher bond dimension and memory. Due to the lower value of the shape parameter $\eta$ we find the bond dimensions required to reach a certain error, however, vary much much more strongly with different covariances matrices than in Fig. 5 where we set $\eta=50$. This is because, for $\eta=1$, matrices are drawn with significant inter-dimensional correlations. These are in general more difficult to represent with a tensor network ansatz due to the presence of significant inter-dimensional correlations between binary digits.
Figure S1: Comparison of the effectiveness of different tree tensor networks (diagrams of the different networks can be found in Fig. 5) with $L=16$ bits per dimension for learning the multinormal probability density function $f(\mathbf{r})\propto\exp(-((\mathbf{r}-\bm{\mu})^{T}M^{-1}(\mathbf{r}-\bm{\mu}))$ via the tensor cross interpolation algorithm. Here $M$ is an $n\times n$ covariance matrix and $\bm{\mu}=(\mu_{1},\mu_{2},...,\mu_{n})$ is the mean vector. We consider $n=3$ with $\mathbf{r}=(x,y,z)\in[0,10)^{3}$ and $\mathbf{\mu}=(5,5,5)$. Results are obtained from drawing $N=10$ instances of $M$ ($M_{1},M_{2},\ldots M_{10}$) from the LKJ distribution [51] with shape parameter $\eta=1$, a distribution uniform in the space of all covariance matrices. a-b) Error $\epsilon$, calculated via Eq. (5) after $n=10$ sweeps of the TCI algorithm, versus bond dimension and memory cost for the tensor networks. The dashed lines shows the mode of the error over the $10$ realizations of $M$ whilst the solid line shows the specific error for $M=M_{1}.$ Inset in a) shows the Kullback-Leibler divergence (calculated from the function samples) as a function of bond dimension evaluated from the function samples for $M=M_{1}$. Inset in b) shows a heatmap of the function for $y=\frac{1}{2}$. c) Average value for the infinity norm $\epsilon_{\infty}$ — see Eq. (5) — over a given sweep of the TCI algorithm for $M=M_{1}$ and the three tensor networks at the specified bond dimensions.

### v2

@@ V Application: Solving Non-linear Fredholm Equations with Tree Tensor Networks
$f(\mathbf{x})=g(\mathbf{x})+\lambda\int_{\mathbf{t}\in[0,1)^{n}}K(\mathbf{x},\mathbf{t})f^{\alpha}(\mathbf{t})d\mathbf{t}\qquad\alpha\in\mathbb{N},$ (7)
@@
  Remap Variables: $f_{i}(\mathbf{t})\leftarrow f_{i}(\mathbf{x})$ $\triangleright$ The mapping is achieved via the index relabelling $x_{i,j}\rightarrow t_{i,j}$.
@@
  Integrate (Sec. V): $f_{i}(\mathbf{x})\leftarrow f_{i}(\mathbf{x},\mathbf{t})\prod_{i=1}^{n}\prod_{j=1}^{L}T_{i,j}(t_{i,j})$ $\triangleright$ $T_{i,j}(t_{i,j})=\frac{1}{2}\ \forall i,j$.
@@
Figure 6: Tree tensor network method for solving (non-linear) Fredholm equations of the second kind $f(\mathbf{x})=g(\mathbf{x})+\int_{\mathbf{t}}f^{\alpha}(\mathbf{t})K(\mathbf{x},\mathbf{t})d\mathbf{t}$. a) An initial guess for the target function $f(\mathbf{x})$ is constructed as a tensor network over a labelled tree. The additive function $g(\mathbf{x})$ is constructed as a tree tensor network of the same structure. The kernel $K(\mathbf{x},\mathbf{t})$ is then constructed over as a tree tensor network formed from two copies of the tree, with one encoding the ‘real’ variables $\mathbf{x}$ and the other the auxiliary variables $\mathbf{t}$. A single edge is also added between the two trees. The Fredholm equation can then be solved by the iterative steps expressed at the bottom and outlined in Alg. 1 b) Results for two different Fredholm equations with a two dimensional solution $f(\mathbf{x})=f(x_{1},x_{2})$. Explicit expressions for $K(x_{1},x_{2},t_{1},t_{2})$ and $g(x_{1},x_{2})$ are given in the main text. The structure of the underlying tree for $f(x)$ is shown at the bottom with the binary digits numbered by their significance: the most significant digit in each dimension is circled in red. Error $\epsilon$ is calculated as in Eq. (6), sampling over 100100 random grid points and comparing to the values for the exact solution, versus iteration of the solver. Dashed lines show the converged error from the Direct Numerical Simulation (DNS) results of Ref. [61] with $N=10,100$ and $1000$ grid points. Left plot) Results for example 1 where the exact solution is $f(x_{1},x_{2})=\sin(x_{2})$. Right plot) Results for example 2 where the exact solution is $f(x_{1},x_{2})=1/(1+x_{1}+x_{2})^{2}$.
@@ Acknowledgements.
We would like to thank Matthew Fishman for helpful discussions and advice and input on software design. We also thank Ryan Anselm and Marc Ritter for useful discussions. The authors are grateful for ongoing support through the Flatiron Institute, a division of the Simons Foundation. The tools and methods described in this paper are contained within the ITensorNumericalAnalysis.jl library [67]: an open-source Julia package built on-top of ITensors.jl [68] and TensorNetworkQuantumSimulator.jl [69] for constructing and optimizing tree tensor network representations of continuous functions, with the tree structure freely definable by the user.
@@ VII Supplementary Material
$=\qquad$ $\mathcal{T}_{(j,j^{\prime})}(x_{j}+x_{j^{\prime}})_{\alpha_{1},\alpha_{2},\ldots,\alpha_{z_{j}-2},\beta}\qquad z_{j}>2,$
@@ VII.3 Further TCI Data
In Fig. S1 we present further data comparing the effectiveness of the tensor cross interpolation (TCI) algorithm for learning the trivariate probability density function $f(\mathbf{r})\propto\exp(-(\mathbf{r}-\mathbf{\mu})^{T}M^{-1}(\mathbf{r}-\mathbf{\mu}))$ where $\mathbf{r}=(x,y,z)$, $\mathbf{\mu}=(\mu_{x},\mu_{y},\mu_{z})$ is the mean vector, and $M$ is a covariance matrix. Here we sample $M$ from the Lewandowski-Kurowicka-Joe (LKJ) [57] distribution with shape parameter $\eta=1$, which is equivalent to sampling uniformly from the space of all covariance matrices. As in the main text, we find the comb tree systematically outperforms the tensor trains for a given covariance matrix $M$. This is most transparent in Fig. S1c), where the TCI algorithm converges to a value for the infinity norm $\epsilon_{\infty}$ that is orders of magnitude lower for the comb tree than for the tensor trains - despite the train ansatzes utilizing a higher bond dimension and memory. Due to the lower value of the shape parameter $\eta$ we find the bond dimensions required to reach a certain error, however, vary much much more strongly with different covariances matrices than in Fig. 5 where we set $\eta=50$. This is because, for $\eta=1$, matrices are drawn with significant inter-dimensional correlations. These are in general more difficult to represent with a tensor network ansatz due to the presence of significant inter-dimensional correlations between binary digits.
Figure S1: Comparison of the effectiveness of different tree tensor networks (diagrams of the different networks can be found in Fig. 5) with $L=16$ bits per dimension for learning the multinormal probability density function $f(\mathbf{r})\propto\exp(-((\mathbf{r}-\mathbf{\mu})^{T}M^{-1}(\mathbf{r}-\mathbf{\mu}))$ via the tensor cross interpolation algorithm. Here $M$ is an $n\times n$ covariance matrix and $\mathbf{\mu}=(\mu_{1},\mu_{2},...,\mu_{n})$ is the mean vector. We consider $n=3$ with $\mathbf{r}=(x,y,z)\in[0,10)^{3}$ and $\mathbf{\mu}=(5,5,5)$. Results are obtained from drawing $N=10$ instances of $M$ ($M_{1},M_{2},\ldots M_{10}$) from the LKJ distribution [57] with shape parameter $\eta=1$, a distribution uniform in the space of all covariance matrices. a-b) Error $\epsilon$, calculated via Eq. (6) after $n=10$ sweeps of the TCI algorithm, versus bond dimension and memory cost for the tensor networks. The dashed lines shows the mode of the error over the 1010 realizations of $M$ whilst the solid line shows the specific error for $M=M_{1}.$ Inset in a) shows the Kullback-Leibler divergence (calculated from the function samples) as a function of bond dimension evaluated from the function samples for $M=M_{1}$. Inset in b) shows a heatmap of the function for $y=\frac{1}{2}$. c) Average value for the infinity norm $\epsilon_{\infty}$ — see Eq. (6) — over a given sweep of the TCI algorithm for $M=M_{1}$ and the three tensor networks at the specified bond dimensions.